Using the following table, please compute f(x.y) 0.s 3.3 a.-0.56 b.0.14 C.0.56 d.-0.14
Answer #1

Lat Total 2. Exo)+ (3x4) = 1.2十1.2十(3x5x0-1) = 10.9-6-3 (2-4) = 109-11-04